OpenStudy (anonymous):

Using the substitution method how to solve the value of x and y. -6x-y=29 / 8x+3y=-47

OpenStudy (saifoo.khan):

So you know how substitution method workd?

OpenStudy (saifoo.khan):

works*

OpenStudy (saifoo.khan):

Try it then! :D

OpenStudy (anonymous):

It's confusing me :(

OpenStudy (saifoo.khan):

Where are you confused?

OpenStudy (anonymous):

on which should i substitute first

OpenStudy (saifoo.khan):

-6x-y=29 -1 8x+3y=-47 -2 -6x - y = 29 Let's solve for y first. y = 29 - 6x Now use this value and substitute them in equation 2.

OpenStudy (saifoo.khan):

8x + 3y = -47 8x + 3(29 - 6x) = -47

OpenStudy (anonymous):

So what should be the right answer?

OpenStudy (saifoo.khan):

Solve this: 8x + 3(29 - 6x) = -47